OpenStudy (rock_mit182):

Two rains leave a train station at the same time. One travels east at 12miles per hour. The other train travels west at 7 miles per hour. In how many hours will the two trains be 114 miles apart

OpenStudy (doodleduck8):

Hi there, Add the two speeds together = 180 Divide distance by 180 216/180 = 1.2 = 1 hour 12 minutes Proof: Eastbound train: Distance = Time x Speed Eastbound = 114 miles West bound train: Distance = Time x Speed Westbound = 102 miles. Added together = 216 miles. Hope this helps :-)

OpenStudy (doodleduck8):

Hi, what do you need help on?

OpenStudy (rock_mit182):

:) THANKS !

OpenStudy (doodleduck8):

No problem!!! :)

OpenStudy (rock_mit182):

As a part of a physics experiment, Ming drops a baseball from the top of a 330-foot building. To the nearest tenth of a second, for how many seconds will the baseball fall? (Hint: use the formula h = 16t^2, which gives the distance h, in feet, that a free-falling object travels in t seconds)

OpenStudy (rock_mit182):

330 = 16 t^2 ?

OpenStudy (rock_mit182):

am i right ?

OpenStudy (doodleduck8):

The given information is that h = 305 We are looking for t h = 16 * t^2 305 = 16 * t^2 19.0625 = t^2 t = 4.4 seconds

OpenStudy (rock_mit182):

to the nearest tenth of a sec that was the missing part.. so i guess i was 90% right

OpenStudy (rock_mit182):

thanks bro!

OpenStudy (doodleduck8):

No problem!! :)